Solve cot θ sin θ = 1/2 for 0 <= θ < 2π.
1
Expert's answer
2012-02-09T08:21:23-0500
cot(theta)*sin(theta) = 1/2.
cot(theta) = cos(theta)/sin(theta), sin(theta)
is not zero. So:
cos(theta)/sin(theta)*sin(theta) = 1/2.
cos(theta) =
1/2.
As theta is from [0,2*pi], then
theta = pi/3 or theta = 5pi/3.